OpenStudy (anonymous):

In a 20 row theater, the number if seats in a row increases by three with each successive row. The first row has 18 seats. a. write an arithmetic series to represent the number if seats in the theater. b. find the total seating capacity of the theater. c. front row tickets for a concert cost $60. After every 5 rows,the tickets price goes down by $5. What is the total amount of money generated by a full house?

OpenStudy (anonymous):

@skullpatrol @ganeshie8 @Jack1

ganeshie8 (ganeshie8):

First row : 18 Second row : 18+3 = ? Third row : 18+3+3 = ? Fourth row : 18+3+3+3 = ?

OpenStudy (anonymous):

so i keep added 3 but i dont get the rest:(

ganeshie8 (ganeshie8):

easy yes :)

ganeshie8 (ganeshie8):

so for part a, total number of seats in the theater is below sum : 18 + 21 + 24 + 27 + ...

ganeshie8 (ganeshie8):

^thats the series. we're done with part a

OpenStudy (anonymous):

thanks what b and c

ganeshie8 (ganeshie8):

for part b, u need to actually add all the seats to get the total number of seats

ganeshie8 (ganeshie8):

remember the partial sum formula for arithmetic series ?

OpenStudy (anonymous):

the 18+3? im lost sorry

ganeshie8 (ganeshie8):

Sum of first "n" terms : \(\large S_n = \dfrac{n}{2}\left[ 2a_1 + (n-1)d\right]\)

OpenStudy (anonymous):

is n 18?

ganeshie8 (ganeshie8):

since there are 20 rows, \(n = 20\) first row has 18 seats, so \(a_1 = 18\) each row is incremented by 3, so \(d = 3\)

ganeshie8 (ganeshie8):

plug the values in formula and simplify

OpenStudy (anonymous):

s= 20/2[2*18 1+(20-1)3] s= 10[36 1+(19) 3]

OpenStudy (anonymous):

s=10[36 (19) 3]

ganeshie8 (ganeshie8):

Yes :) \(\large S_n = \dfrac{n}{2}\left[ 2a_1 + (n-1)d\right]\) \(\large S_{20} = \dfrac{20}{2}\left[ 2*18 + (20-1)*3\right]\) \(\large ~~~~~= 10\left[ 36 + (19)*3\right]\)

ganeshie8 (ganeshie8):

^^simplify

OpenStudy (anonymous):

10(36+19 times 3) ?

OpenStudy (anonymous):

=930

ganeshie8 (ganeshie8):

Perfect ! so, the total seating capacity of theater is \(930\)

OpenStudy (anonymous):

thanks so what about c

ganeshie8 (ganeshie8):

for part C, u need to use the same formula FOUR times

ganeshie8 (ganeshie8):

c. front row tickets for a concert cost $60. After every 5 rows,the tickets price goes down by $5. What is the total amount of money generated by a full house?

ganeshie8 (ganeshie8):

Since there are 20 rows in the theater, Cost of sitting in First 5 rows(1st row to 5th row) = 60 Cost of sitting in next 5 rows(6th row to 10th row) = 55 Cost of sitting in next 5 rows(11th row to 15th row) = 50 Cost of sitting in next 5 rows(15th row to 20th row) = 45

OpenStudy (anonymous):

5th 60 10th 55 15 50 20 45

ganeshie8 (ganeshie8):

Exactly ! To get the money generated in First 5 rows : tyou need to find out how many total seats are there in first 5 rows, and multiply it by 60

ganeshie8 (ganeshie8):

use the same formula, just replace \(n\) by \(5\)

ganeshie8 (ganeshie8):

\(\large S_n = \dfrac{n}{2}\left[ 2a_1 + (n-1)d\right] \) \(\large S_5 = ? \)

OpenStudy (anonymous):

s5= 5/2 [2*18+(5-1)3} s5=2.5(26+4 times 3)

ganeshie8 (ganeshie8):

simplify all the way

ganeshie8 (ganeshie8):

and there is a mistake in ur second step : 2*18 = 36, NOT 26 okay ?

OpenStudy (anonymous):

750/20=37.5

ganeshie8 (ganeshie8):

No wait, from where u getting these numbers lol

OpenStudy (anonymous):

2.5 (120) 15 seats

ganeshie8 (ganeshie8):

\(\large S_n = \dfrac{n}{2}\left[ 2a_1 + (n-1)d\right] \) \(\large S_5 = \dfrac{5}{2}\left[ 2*18 + (5-1)*3\right] \) \(\large ~~~= \dfrac{5}{2}\left[ 36+ 12\right] \) \(\large ~~~= \dfrac{5}{2}\left[ 48\right] \) \(\large ~~~= 5\left[ 24\right] \) \(\large ~~~= 120 \)

ganeshie8 (ganeshie8):

^^ So, there are aotal of t\(120\) seats in the first 5 rows.

ganeshie8 (ganeshie8):

Multiply it by 60 to get the money earned

OpenStudy (anonymous):

so 6 seats per row?

OpenStudy (anonymous):

so 360

ganeshie8 (ganeshie8):

Nope. there are a total of \(120\) seats in the first 5 rows. and each seat earns 60. So total money earned on first 5 rows = \(120 \times 60 = ?\)

OpenStudy (anonymous):

7200?

ganeshie8 (ganeshie8):

Yep !

ganeshie8 (ganeshie8):

`120 sears` in the first 5 rows earn a money of `7200`

ganeshie8 (ganeshie8):

*120 seats

ganeshie8 (ganeshie8):

Okay ? similarly u need to calculate how many seats are there in next 5 rows and multiply it by 55

OpenStudy (anonymous):

thanks so much that it thanks again

OpenStudy (anonymous):

lol i though were done haha

ganeshie8 (ganeshie8):

np :) but the part c is not over yet... there are 15 more rows in the back lol

ganeshie8 (ganeshie8):

u dont want more money ha ?

ganeshie8 (ganeshie8):

u just calculated the earning on first 5 rows.

OpenStudy (anonymous):

lol im tried

ganeshie8 (ganeshie8):

15 more rows are yet to be calculated lol

OpenStudy (anonymous):

haha yea ;)

ganeshie8 (ganeshie8):

Number of seats in next 5 rows (6th row to 10th row) : http://www.wolframalpha.com/input/?i=10%2F2%282*18%2B%2810-1%29*3%29+-+120

ganeshie8 (ganeshie8):

multiply it by 55

OpenStudy (anonymous):

195 times 55 =10725

ganeshie8 (ganeshie8):

Wow ! thats lots of money !!

ganeshie8 (ganeshie8):

Add it to ur previous amount

ganeshie8 (ganeshie8):

your first 5 rows earned `7200` ?

ganeshie8 (ganeshie8):

So total money earned from 1st row to 10th row = `7200+10725` = ?

OpenStudy (anonymous):

lol someone is getting rich haha

ganeshie8 (ganeshie8):

Party !!!!

OpenStudy (anonymous):

17925 lol why not party all day long haha

ganeshie8 (ganeshie8):

:) lets get greedy... see how much next 5 rows earn u : total number of seats in next 5 rows (11th row to 15th row) : http://www.wolframalpha.com/input/?i=15%2F2%282*18%2B%2815-1%29*3%29+-+%28120%2B195%29

ganeshie8 (ganeshie8):

multiply it by 50

OpenStudy (anonymous):

lol but i want to get a audi first haha

OpenStudy (anonymous):

13500

ganeshie8 (ganeshie8):

hahah sure why not ;)

OpenStudy (anonymous):

what about u what kind of car u want

ganeshie8 (ganeshie8):

yes add that to ur previous money : `7200+10725 + 13500`

ganeshie8 (ganeshie8):

^^thats the money earned by ur first 15 rows

OpenStudy (anonymous):

31425

ganeshie8 (ganeshie8):

yes ! this is the number of seats in last 5 rows : http://www.wolframalpha.com/input/?i=20%2F2%282*18%2B%2820-1%29*3%29+-+%28120%2B195%2B270%29

ganeshie8 (ganeshie8):

Multiply it by 45 and Enjoy.... !!

OpenStudy (anonymous):

15525

OpenStudy (anonymous):

46950 is the final answer

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

here is ur gift for getting the correct answer everytime http://choicewallpapers.net/wp-content/uploads/2013/05/Audi-car6.jpg

ganeshie8 (ganeshie8):

gtg... cya :)

OpenStudy (anonymous):

thank you so much thanks for helping me understand this problem

OpenStudy (anonymous):

thanks i love that car its perfect